# Quantitative agreement between experiment and theory for Vibrational Circular Dichroism enhanced by electronically excited states

### v1

@@ 5.1 Magnetic field perturbation calculations
Zn(II) has a closed-shell $d^{10}$ configuration with high-lying excited states, so enhancement effects for Zn(sp)Cl2 are negligible and the MFP approximation is fully justified. As a result, this complex serves as a good benchmark for MFP-based VCD calculations. The IR spectra of all three studied M(sp)Cl2 (M = Zn, Co, Ni) are almost identical, so the Zn(sp)Cl2 complex can be used to select the most appropriate DFT functional and to evaluate conformational effects. Conformational analysis reveals an energy difference between the two lowest energy conformers of Zn(sp)Cl2 of 4.2 kcal/mol which exceeds the value 3 kcal/mol typically used to select contributing conformers. This is in agreement with previous studies of the sparteine ligand  35 who also found Me(II)(sp)Cl2 complexes to exist predominantly in a single conformation and makes the systems ideal for analysing their VCD spectra in detail.
@@
Several benchmark studies have shown that the generalized gradient approximation (GGA) functional BP86 39, 52 outperforms B3LYP 53 in reproducing experimental equilibrium geometries of small first-row transition metal complexes 54, 55. Minenkov et al.56 furthermore reported that B3LYP systematically overestimates the metal-ligand bonds in transition metal complexes of the first and second row. A more recent benchmark for metal–dinitrogen complexes 57 further showed that DFT performance is highly sensitive to the nature of the N2 coordination mode at the metal center. We compare the performance of three DFT functionals (BP86, B3LYP, PBE0 58 with and without Grimme’s D3(BJ) dispersion correction  41, 42 in the gas phase and with CPCM of CHCl3  43 (see detailed plots in the Supporting Information).
@@
Figure 2 shows IR and MFP-VCD of Zn(sp)Cl2 with the considered functionals. The addition of the dispersion correction improves both simIR and simVCD for BP86 and B3LYP (see Figure S1 in SI). BP86-D3BJ/def2-TZVP with CPCM demonstrates the best agreement with the experiment with good similarities for both IR and VCD (simIR = 0.88, simVCD = 0.44). Typically, simVCD values larger than 0.4 are considered to be secure to assign the absolute configuration 59. The calculated IR spectra of open-shell Co(sp)Cl2 and Ni(sp)Cl2Figure 3) at the same level of theory have a high similarity to the experiment. This confirms the suitability of BP86-D3BJ/def2-TZVP with CPCM to optimize the ground state structure and compute IR spectra for these compunds.
@@
Figure 2: Comparison of VCD (top) and IR (bottom) spectra of Zn(sp)Cl2. Experimental spectra are given in black, the spectra are calculated with CPCM of CHCl3 with B3LYP/def2-TZVP (red, simVCD = 0.16), PBE0/def2-TZVP (green, simVCD = 0.28), BP86/def2-TZVP (blue, simVCD = 0.44).
@@
Figure 3: Co(sp)Cl2 calculated (blue, simIR = 0.88) and observed (black) IR spectra (left). Ni(sp)Cl2 calculated (red, simIR = 0.85) and observed (black) IR spectra (right). All calculations are performed with BP86-D3BJ/def2-TZVP with CPCM of CHCl3.
@@ 5.2 Excited state calculations
We next investigate the excitation energies and the nature of LLES for the Co(II)(sp)Cl2 and Ni(II)(sp)Cl2 complexes with TDDFT and SA-CASSCF. The contributions of each excited state to the enhanced VCD spectra depend on the EDTM, MDTM, excitation energy, nonadiabatic coupling matrix elements, and vibrational frequencies, so the quality of all these quantities is important to obtain a reliable estimate of the enhancement that can be expected in VCD.
According to experimental studies, pseudotetrahedral M(sp)Cl2 complexes exhibit low-lying, dipole-forbidden $d-d$ transitions that underlie their enhancement effect 21. Our SA-CASSCF calculations on Co(II)(sp)Cl2 and Ni(II)(sp)Cl2 confirm the dominant $d$ character of active orbitals with Löwdin orbital compositions 94% and higher (see Figure 4). In the main text, we present results for minimal active spaces (7,5) for Co(II) and (8,5) for Ni(II), while data for the larger active spaces for Co(sp)Cl2 is provided in the Supporting Information (Table S2). Since the transitions of interest are highly localized, we also construct a smaller model complex, M(II)(NH3)2Cl2, with $C_{2v}$ symmetry. This higher symmetry allows state specific CASSCF calculations, thereby avoiding possible artifacts introduced by orbital averaging in SA-CASSCF.
@@
Co(sp)Cl2 SA-CASSCF(7, 5) calculations show that the first three excited states lie approximately 0.2 eV below the next higher state (see  Figure 5). Since each state’s contribution scales inversely with $w_{eg}-w_{a}$, only those excited states whose energies nearly coincide with a vibrational mode yield significant enhancement. Consequently, contributions from the fourth and higher excited states are effectively negligible. For Ni(sp)Cl2, there is already a significant gap between the first and second excited states (see  Figure 5) and for this complex only the first excited state needs to be considered in the enhancement calculations. In  Figure 5, we compare the excitation energies computed with representative GGA, hybrid, and hybrid meta-GGA functionals. Pure GGAs substantially overestimate the energies of the low-lying states for a high-spin complex. Introducing Hartree–Fock exchange systematically improves the agreement: B3LYP (20% HF exchange) reduces the error, and BHandH (50% HF exchange) performs even better. In contrast, the hybrid meta-GGA M06 60 yields a too low first excitation energy relative to the SA-CASSCF benchmark values.
@@
Figure 5: Excitation energies for the first 6 excited states calculated with TDDFT and SA-CASSCF in Orca 44 for Co(sp)Cl2 (left) and Ni(sp)Cl2 (right). We use the symmetry labels from C2v model (see Table 1 complex to distinguish the excited states for non-symmetric sparteine. The functionals are ordered from GGA (BP86) to hybrid functionals with increasing Hartree-Fock exchange: B3LYP (20%), PBE0 (20%), BHandH (50%). The Hybrid Meta-GGA functional M06 yields a too low first excitation energy, and is put to the right of the reference results to ease the comparison.
@@
Table 1 summarizes CASSCF/NEVPT2 excitation energies for the sparteine and symmetric model complexes. Notably, the SA-CASSCF(7,5) results for Co(sp)Cl2 and Co(NH3)2Cl2 are nearly identical, confirming the highly localized character of these transitions. For the $C_{2v}$ complex the ground (${}^{4}A_{2}$) and excited states (${}^{4}A_{1}$, ${}^{4}B_{1}$ and ${}^{4}B_{2}$) each belong to different irreducible representations, which permits state specific CASSCF optimizations. Here, the first two excited states are almost degenerate, while the third lies slightly higher in energy. For Ni(II)(NH3)3Cl2 complex, the ground state is ${}^{3}B_{2}$, and its three lowest excited states are ${}^{3}A_{2}$, ${}^{3}B_{1}$, and ${}^{3}B_{2}$. Overall, the close agreement between the SS-CASSCF energies of the high-symmetry model and the SA-CASSCF results confirms that the state-averaged orbitals provide a reliable description of these local excitations.
@@
Table 1: Vertical (v) and adiabatic (a) excitation energies of sparteine and model complexes calculated with CASSCF method. NEVPT2 corrected energies are given in brackets. All excitation energies are given in eV.
@@
One key approximation in the Nafie framework that we already mentioned when discussing his theory is the vertical approximation. To check how this approximation can affect results, we show in the last column of Table 1 the adiabatic excitation energies of the sparteine complexes. For the ground state we retain the BP86 geometry, as it yields geometries in best agreement with experimental data (see Figure S2 in SI). For the excited states, however, we decided to optimize geometries using the BHandH functional which provides a state description more closely resembling that obtained from SA-CASSCF than is obtained with TDDFT employing BP86 (see additional results in Table S3 in SI). We thereby assume that, since the calculation of the TDDFT excited-state gradients requires evaluation of many terms shared with a non-adiabatic couplings (NACs) calculation within the linear response framework, a functional that yields more accurate NACs is likely to provide a more reliable description of excited-state gradients.
In Co(sp)Cl2, excited state relaxation lowers the excitation energy by more than 0.10 eV, which is significant in case of such small energies. This shows the limitation of using a vertical approximation, and indicates that inclusion of Franck-Condon factors is likely necessary to achieve quantitatively accurate enhanced VCD predictions by fully ab initio calculation. As this will lead to impractically large computational demands, we will in the next sections demonstrate that a semi-empirical approach can be used instead.
5.3 SA-CASSCF calculations of the enhancement
@@
We compute the enhanced VCD spectra using our Python implementation of Equation 16. The APT and AAT tensors are evaluated at the BP86-D3BJ/CPCM level, and corrections to these tensors are obtained from SA-CASSCF calculations in OpenMolcas averaged over 4 states. The results of VCD calculations for Co(sp)Cl2 are given in Figure 6. Incorporation of three excited states with unmodified vertical excitation energies produces a slight increase in simVCD compared to the MFP, but fails to reproduce the pronounced amplification observed experimentally. Given the uncertainty due to the accuracy of the computed excitation energy and the effect of the vertical approximation, we decided therefore to treat excited state energies as tunable parameters which are optimized to maximize the simVCD value. This is inspired by a similar approach used by some of us in earlier work to deal with uncertainties in determining Boltzmann weights in conformational averaging59. In our current approach the transition dipole moments and non-adiabatic couplings are all unchanged and taken directly from the SA-CASSCF calculations, having only the three excited state energies as tunable parameters. Using this approach we can achieve reliable agreement with the experimental spectra with a simVCD value of 0.69. The energies of the first two excited states converge to an identical value of 0.23 eV, which is approximately 0.1 eV lower than the SA-CASSCF value. This deviation is well within the expected error of this method. The energy of the third state did not change by optimization and remained at 0.42 eV, thereby reducing the number of required tunable parameters required to just two.
@@
Figure 6: Enhanced VCD spectra of Co(sp)Cl2 calculated using MFP (red, simVCD = 0.15), SA-CASSCF (green, simVCD = 0.30), and optimized excited state energies (blue, simVCD = 0.69) are demonstrated on the left. Dependence of the VCD intensity on the energy value of the first two excited states is shown on the right.
@@
Figure 6 illustrates the sensitivity of the simulated VCD enhancement to the energies of the first two excited states. The calculated peaks are slightly shifted relative to the experimental spectrum because we applied only a single frequency scaling factor from the IR calculations. Even 0.02 eV changes in excitation energies significantly affects the amplification of the 1300–1500 cm-1 bands. Such small energy difference far exceeds the typical uncertainty of even high-level quantum chemistry methods. Consequently, fitting low-lying excitation energies to experimental VCD data would be practical route to reproduce the observed enhancement.
@@
Figure 7: Individual contributions of the excited states to the enhancement for Co(sp)Cl2 with the optimized energies. The first two states have both excitation energies 0.23 eV and give similarity 0.38 (dark blue) and 0.29 (blue) respectively. The excitation energy of the third state is 0.42 eV and it’s contribution almost zero (light blue). The grey line corresponds to the excitation energy of the third state shifted to 0.23 eV, the simVCD is -0.11 eV in this case.
@@
The individual contributions of each excited state to the enhanced VCD spectrum are shown in  Figure 7. Considering only the first or only the second excited states increases simVCD values by 0.38 and 0.29, respectively. In contrast, employing the third excited state (at 0.42 eV) gives virtually no similarity to the observed spectrum (simVCD = –0.01, light-blue trace), indicating that its energy lies above the resonance window required for enhancement. When the third state’s energy is artificially shifted downward to 0.23 eV (gray trace), its simVCD even has an inverted sign, giving further confirmation that it does not contribute to the experimentally observed spectrum. From these results, we conclude that only the two lowest excited states drive the vibronic enhancement in Co(sp)Cl2.
@@
For Ni(sp)Cl2 the enhanced spectra with unmodified SA-CASSCF energies also show only slight enhancement compared to MFP spectra  Figure 8). Optimization of the excitation energies including 3 lowest excited states gives a simVCD value of 0.51 with tuned energies of 0.24, 0.31, and 0.37 eV. Including only the lowest state in the optimization yields virtually identical spectra, indicating that in this case only one state contributes to the enhancement. This observation is consistent with the calculated SA-CASSCF excitation energies which predict the second and third excited states to lie significantly higher than the first. After the optimization the calculated Ni(sp)Cl2 spectrum shows one important difference with the experimentally observed one: in the calculated enhanced spectra an intense negative peak appears around 1300 cm-1. This discrepancy already appears in the MFP spectra and may for instance arise from artifacts due to the implicit modeling of solvent effects. If we calculate the spectra similarity excluding this 1300-1320 cm-1 region, where the discrepancy is observed the simVCD is increased to 0.57.
@@
We conclude that SA-CASSCF, with tuning of the lowest excitation energies, can yield very reasonable enhanced VCD spectra for the sparteine complexes. However, SA-CASSCF calculations of non-adiabatic couplings will present a serious bottleneck for applying this approach to larger systems. To be able to overcome this computational bottleneck, we will next explore TDDFT as a more scalable alternative to model VCD intensity enhancement.
5.4 TDDFT calculations of the enhancement
Most density functional approximations that we studied predict excitation energies that are too high to contribute to the enhancement (see additional results in the Table S4 and S5 in SI). To remedy this method artifact, we optimize the excitation energies in the same way as in the SA-CASSCF approach, while retaining the transition dipole moments and non-adiabatic couplings from the TDDFT calculations. BHandH and B3LYP yield nearly identical simVCD values; here, we report the BHandH results since they give the largest simVCD (see  Figure 9). Data for B3LYP and other DFAs is available in the Supporting Information.
Figure 9: Enhanced VCD spectra of Co(sp)Cl2 calculated with BHandH functional including two excited states (red, simVCD = 0.51) and three excited states (blue, simVC = 0.67) on the left. Individual contributions of the first three excited states to the enhancement for Co(sp)Cl2 from BHandH calculation with similarities 0.20, 0.25 and 0.16.
For the Co(sp)Cl2 complex, optimization of three excited states energies produces close values around 0.23 eV. This differs from the SA-CASSCF results, where the third state remained higher and did not contribute to the enhancement. We link such differences to the non-adiabatic couplings (NACs) (see Figure 10). The only significant coupling values correspond to atoms directly bonded to the transition metal, again confirming the distance-dependence of the enhancement effect. The NACs between the ground and second excited states are very close for BHandH and SA-CASSCF. More significant deviations are observed for the first and the third state. In the third state, the couplings involving the N atoms are noticeably stronger with BHandH, which likely enables it’s contribution to the enhancement. We find that the non-adiabatic coupling vectors computed with B3LYP closely match those obtained with the BHandH DFA, which accounts for the nearly identical enhanced VCD profiles predicted by both methods. NACs, difference plots and mean errors for other DFAs can be found in the SI.
(a) NACs between the ground and first excited state
(b) NACs between the ground and second excited state
(c) NACs between the ground and third excited state

### v2

@@ 5.1 Magnetic field perturbation calculations
Zn(II) has a closed-shell $d^{10}$ configuration with high-lying excited states. Enhancement effects are in this case thus negligible and the MFP approximation is fully justified. As a result, this complex serves as a good benchmark for MFP-based VCD calculations. The IR spectra of all three studied M(sp)Cl2 (M = Zn, Co, Ni) are almost identical, so the Zn(sp)Cl2 complex can be used to select the most appropriate DFT functional and to evaluate conformational effects. Conformational analysis reveals an energy difference between the two lowest energy conformers of Zn(sp)Cl2 of 4.2 kcal/mol which exceeds the value of 3 kcal/mol typically used to select contributing conformers. This is in agreement with previous studies of the sparteine ligand  35 where it was also found that Me(II)(sp)Cl2 complexes exist predominantly in a single conformation, which makes the systems ideal for analysing their VCD spectra in detail.
@@
Several benchmark studies have shown that the generalized gradient approximation (GGA) functional BP86 39, 52 outperforms B3LYP 53 in reproducing experimental equilibrium geometries of small first-row transition metal complexes 54, 55. Minenkov et al.56 furthermore reported that B3LYP systematically overestimates the metal-ligand bonds in transition metal complexes of the first and second row. A more recent benchmark for metal–dinitrogen complexes 57 further showed that the performance of DFT is highly sensitive to the nature of the N2 coordination mode at the metal center. In view of these observations and to establish the optimal computational approach for the present system we compare in the following the performance of three DFT functionals (BP86, B3LYP, PBE0) 58 with and without Grimme’s D3(BJ) dispersion correction  41, 42 in the gas phase and with CPCM of CHCl3  43 (see detailed plots in the SI).
@@
Figure 2 shows calculated IR and MFP-VCD spectra of Zn(sp)Cl2 with the considered functionals. The addition of the dispersion correction improves both simIR and simVCD for BP86 and B3LYP (see Figure S1 in SI). BP86-D3BJ/def2-TZVP with CPCM demonstrates the best agreement with the experiment with good similarities for both IR and VCD (simIR = 0.88, simVCD = 0.44), noticing that simVCD values larger than 0.4 are typically considered to be secure to assign the absolute configuration 59.  Figure 3 shows that this level of theory predicts IR spectra for the open-shell Co(sp)Cl2 and Ni(sp)Cl2 systems that also have a high similarity to the experiment. We thus conclude that calculations at the BP86-D3BJ/def2-TZVP with CPCM level of theory are suitable to optimize the ground state structure of these compoundsand compute their IR spectra .
@@
Figure 2: Comparison of VCD (top) and IR (bottom) spectra of Zn(sp)Cl2. Experimental spectra are given in black, the spectra are calculated with CPCM of CHCl3 with PBE0/def2-TZVP (red, simVCD = 0.16), B3LYP/def2-TZVP (green, simVCD = 0.28), and BP86/def2-TZVP (blue, simVCD = 0.44).
@@
Figure 3: Left panel: calculated (blue, simIR = 0.88) and observed (black) IR spectra of Co(sp)Cl2. Right panel: calculated (red, simIR = 0.85) and observed (black) IR spectra of Ni(sp)Cl2. All calculations have been performed with BP86-D3BJ/def2-TZVP with CPCM of CHCl3.
@@ 5.2 Excited state calculations
We next investigate with TDDFT and SA-CASSCF the excitation energies and nature of LLES for the Co(II)(sp)Cl2 and Ni(II)(sp)Cl2 complexes. The contributions of each excited state to the enhanced VCD spectra depend on the EDTM, MDTM, excitation energy, nonadiabatic coupling matrix elements, and vibrational frequencies, so the quality of all these quantities is important to obtain a reliable estimate of the enhancement that can be expected in VCD.
According to experimental studies, pseudotetrahedral M(sp)Cl2 complexes exhibit low-lying, dipole-forbidden $d$-$d$ transitions that underlie their enhancement effect 21. Our SA-CASSCF calculations on Co(II)(sp)Cl2 and Ni(II)(sp)Cl2 confirm the dominant $d$ character of active orbitals with Löwdin orbital compositions 94% and higher (see Figure 4). In the main text, we present results for minimal active spaces of (7,5) for Co(II) and (8,5) for Ni(II), while data for the larger active spaces for Co(sp)Cl2 is provided in the SI (Table S2). Since the transitions of interest are highly localized, we also construct a smaller model complex (M(II)(NH3)2Cl2) with $C_{2v}$ symmetry. This higher symmetry allows state-specific CASSCF calculations, thereby avoiding possible artifacts introduced by orbital averaging in SA-CASSCF.
@@
Co(sp)Cl2 SA-CASSCF(7, 5) calculations show that the first three excited states lie approximately 0.2 eV below the next higher state (see  Figure 5). Since the contribution of each state scales inversely with $w_{eg}-w_{a}$, only excited states with energies that nearly coincide with the frequency of a vibrational mode yield significant enhancements. Contributions from the fourth and higher excited states are thus for all practical purposes negligible. For Ni(sp)Cl2 we find on the other hand that there is already a significant energy gap between the first and second excited state (see  Figure 5), and for this complex we therefore only need to consider the first excited state in the enhancement calculations.
In order to assess to what extent excitation energies calculated at the TDDFT level are able to reproduce these SA-CASSCF energies, we compare in  Figure 5 the latter energies with excitation energies computed with representative GGA, hybrid, and hybrid meta-GGA functionals. From this Figure it becomes clear that pure GGAs substantially overestimate the energies of the low-lying states for a high-spin complex. Introducing Hartree–Fock exchange systematically improves the agreement: B3LYP (20% HF exchange) reduces the error, and BHandH (50% HF exchange) performs even better. In contrast, the hybrid meta-GGA M06 60 yields a first excitation energy that is too low relative to the SA-CASSCF benchmark values.
@@
Figure 5: Excitation energies of the first 6 excited states of Co(sp)Cl2 (left) and Ni(sp)Cl2 (right) calculated with TDDFT and SA-CASSCF in Orca 44. To distinguish the excited states we use for non-symmetric sparteine the symmetry labels of the $C_{2v}$ model complex (see Table 1. The functionals are ordered from GGA (BP86) to hybrid functionals with increasing Hartree-Fock exchange: B3LYP (20%), PBE0 (20%), BHandH (50%). The Hybrid Meta-GGA functional M06 yields a first excitation energy that is too low, and is put to the right of the reference results to facilitate the comparison.
@@
In order to determine whether the use of state-averaged orbitals for the sparteine complexes does not introduce mixing of the character of the electronic states, we compare in Table 1 CASSCF/NEVPT2 excitation energies for the sparteine and symmetric model complexes. Notably, the SA-CASSCF(7,5) results for Co(sp)Cl2 and Co(NH3)2Cl2 are nearly identical, confirming the highly localized character of these transitions. For the $C_{2v}$ complex the ground (${}^{4}A_{2}$) and excited states (${}^{4}A_{1}$, ${}^{4}B_{1}$ and ${}^{4}B_{2}$) belong each to different irreducible representations, which permits state-specific CASSCF optimizations. Here, the first two excited states are almost degenerate, while the third state lies slightly higher in energy. For the Ni(II)(NH3)3Cl2 complex, the ground state is ${}^{3}B_{2}$, while the three lowest excited states are the ${}^{3}A_{2}$, ${}^{3}B_{1}$, and ${}^{3}B_{2}$ states. Overall, the close agreement between the SS-CASSCF energies of the high-symmetry model and the SA-CASSCF results confirms that the state-averaged orbitals provide a reliable description of these local excitations.
@@
Table 1: Vertical (v) and adiabatic (a) excitation energies (eV) of sparteine and model complexes calculated at the CASSCF level. NEVPT2 corrected energies are given in brackets.
@@
As discussed above, one of the key approximations in the SOS formalism is that the potentials energy surfaces of ground and excited states are not displaced with respect to each other. To check whether this assumption holds for the present systems, we show in the last column of Table 1 the adiabatic excitation energies of the Co(sp)Cl2 and Ni(sp)Cl2 complexes. For the ground state we retain the BP86 geometry as it yields geometries in best agreement with experimental data (see Figure S2 in SI). For the excited states, however, we decided to optimize geometries using the BHandH functional, because it provides a state description that resembles more closely the one obtained from SA-CASSCF than the description obtained with TDDFT employing BP86 (see additional results in Table S3 in SI). We thereby assume that – since the calculation of the TDDFT excited-state gradients requires evaluation of many terms shared with a non-adiabatic couplings (NACs) calculation within the linear response framework – a functional that yields more accurate NACs is likely to provide a more reliable description of excited-state gradients. For Co(sp)Cl2 we find that geometry relaxation of the first excited state relaxation lowers the excitation energy by more than 0.10 eV, which is more than a third of the vertical excitation energy and thus quite significant. For the second excited state, on the other hand, the relaxation energy is much smaller though still more than 10% of its vertical excitation energy. Finally, for the third excited state a significant relaxation energy is observed as well. For Ni(sp)Cl2 similar observations are made although in this case the relaxation energy of the first excited state is somewhat lower.
The important conclusion that we thus need to draw is that applying the vertical approximation within the SOS formalism will have a significant impact on the outcome of the calculations. [^1] In passing by, we note that a similar assumption is made as well in the MFP formalism but in that case its repercussions can be expected to be much less. To achieve quantitatively accurate enhanced VCD predictions by fully ab initio calculations would require explicit calculations of vibrational overlap integrals and their inclusion in the calculation. This will lead, however, to impractically large computational demands, and alternative solutions are needed. In the next sections we will therefore introduce a readily usable semi-empirical approach and demonstrate that such an approach allows for calculations that quantitatively reproduce experimental VCD spectra.
5.3 SA-CASSCF calculations of enhancements
@@
We compute enhanced VCD spectra using our Python implementation of Equation 16. To this purpose the APT and AAT tensors are evaluated at the BP86-D3BJ/CPCM level, and corrections to these tensors are obtained from OpenMolcas SA-CASSCF calculations averaged over 4 states. The results of VCD calculations for Co(sp)Cl2 are shown in the left panel of Figure 6. Inclusion of three excited states with unmodified vertical excitation energies produces a slight increase in simVCD compared to the MFP calculation, but fails to reproduce the pronounced amplification observed experimentally. Given the uncertainty due to the accuracy of the computed excitation energies and the previously discussed caution that needs to be taken in using the vertical approximation, we decided to treat the excited state energies as tunable parameters and optimize them as to maximize the simVCD value. Such an approach closely follows the approach that we used in earlier work to deal with uncertainties in determining Boltzmann weights in conformational averaging59. We thus take the transition dipole moments and non-adiabatic couplings directly from the SA-CASSCF calculations and leave them unchanged, but use the excited-state energies of the three lower excited states as tunable parameters. Using this approach we can achieve excellent agreement with the experimental spectrum of Co(sp)Cl2 with a simVCD value of 0.69. The energies of the first two excited states converge to a value of 0.23 eV, which is approximately 0.1 eV lower than the value obtained with SA-CASSCF calculations. Such a difference is quite acceptable and is well within the expected error of this method. We find that the energy of the third state does not change by optimization and remains at 0.42 eV, thereby reducing the number of tunable parameters required to just two.
@@
Figure 6: Left panel: enhanced VCD spectra of Co(sp)Cl2 calculated using MFP (red, simVCD = 0.15), SA-CASSCF (green, simVCD = 0.30), and optimized excited state energies (blue, simVCD = 0.69). Right panel: dependence of the VCD intensity in the 1430-1500 cm-1 region on the excitation energies of the first two excited states.
@@
The right panel in  Figure 6 illustrates the sensitivity of the simulated VCD enhancement to the energies of the first two excited states. The calculated bands are slightly shifted relative to the experimental spectrum because we applied only a single frequency scaling factor from the IR calculations. More important is that the Figure convincingly demonstrates that even a minor change of 0.02 eV in excitation energy affects dramatically the amplification of the 1430–1500 cm-1 bands. Such a small energy difference far exceeds the typical uncertainty of even the highest-level quantum chemistry methods available. Consequently, fitting low-lying excitation energies to experimental VCD data would be more than justified and is a quite practical route to reproduce the observed enhancements.
@@
Figure 7: Individual contributions of excited states with optimized energies to the enhancement of VCD bands in the VCD spectrum of Co(sp)Cl2. The first two states both have excitation energies 0.23 eV and give rise to a simVCD value of 0.38 (dark blue) and 0.29 (blue), respectively. The excitation energy of the third state is 0.42 eV and its contribution is almost zero (light blue). The grey line corresponds to the excitation energy of the third state shifted to 0.23 eV in which case a simVCD value of -0.11 is obtained.
@@
The individual contributions of each excited state to the enhanced VCD spectrum are shown in  Figure 7. Considering only the first or second excited state increases the simVCD value to 0.38 and 0.29, respectively. In contrast, using the third excited state, which is at 0.42 eV, gives virtually no similarity to the observed spectrum (simVCD = –0.01, light-blue trace), indicating that its energy lies above the resonance window required for enhancement. When the energy of the third state is artificially shifted to 0.23 eV (gray trace) its simVCD even has an inverted sign, giving further confirmation that it does not contribute to the experimentally observed spectrum. From these results, we conclude that only the two lowest excited states are responsible for the vibronic enhancement in Co(sp)Cl2. One final aspect that is important to recognize is that the individual contributions show that each state contributes differently both in sign and amplitude to the enhancement of the various bands in the VCD spectrum. The enhancements observed in the experimental spectrum thus actually provide a fingerprint of the extent to which each excited state is of importance in determining the enhancement of that band.
@@
For Ni(sp)Cl2 the enhanced spectra with unmodified SA-CASSCF energies also show only a slight enhancement compared to MFP spectra  Figure 8). Optimization of the excitation energies of the 3 lowest excited states gives a simVCD value of 0.51 with optimized energies of 0.24, 0.31, and 0.37 eV. Including only the lowest state in the optimization yields virtually identical spectra, indicating that in this case only one state contributes to the enhancement. This observation is consistent with the calculated SA-CASSCF excitation energies which predict that the second and third excited states lie significantly higher than the first one. After the optimization the calculated Ni(sp)Cl2 spectrum shows one noticable difference with the experimentally observed spectrum in that in the calculated enhanced spectra an intense negative peak appears around 1300 cm-1. This discrepancy already appears in the MFP spectra and might be due to artifacts arising from the implicit modeling of solvent effects. If we exclude this 1300-1320 cm-1 region, a simVCD value of 0.57 is obtained. Irrespective of whether this region is included or not, the overall conclusion remains, however, that our approach enables a reliable assignment of the absolute configuration of the complex.
@@
We conclude that by tuning the lowest excitation energies, SA-CASSCF is able to predict VCD spectra of sparteine complexes that are in excellent agreement with the experimental spectra. However, SA-CASSCF calculations of non-adiabatic couplings will present a serious bottleneck for applying this approach to larger systems. To overcome this computational bottleneck, we will in the next section demonstrate that TDDFT can fruitfully be employed as a more scalable alternative to model VCD intensity enhancements.
5.4 TDDFT calculations of enhancements
Most density functional approximations that we studied predict excitation energies that are too high to contribute to the enhancements (see additional results in Table S4 and S5 in SI). To remedy this, we optimize the excitation energies in the same way as in the SA-CASSCF approach, while retaining the transition dipole moments and non-adiabatic couplings from the TDDFT calculations. We find that calculations using BHandH and B3LYP yield nearly identical simVCD values, and we therefore report here the BHandH results (see  Figure 9). Data for B3LYP and other DFAs is available in the SI.
Figure 9: Left panel: Enhanced VCD spectra of Co(sp)Cl2 calculated with the BHandH functional including two (red, simVCD = 0.51) and three excited states (blue, simVC = 0.67). Right panel: Individual contributions of the first three excited states to the enhancements in the VCD spectrum of Co(sp)Cl2 using BHandH leading to simVCD values of 0.20, 0.25 and 0.16, respectively.
For the Co(sp)Cl2 complex, optimization of the excitation energies of the three lowest excited states leads for all three states to an excitation energy around 0.23 eV. This differs from the SA-CASSCF results, where the third state remains higher and does not contribute to the enhancement. We attribute these differences to differences in non-adiabatic couplings (NACs) (see Figure 10). The only significant coupling values correspond to atoms directly bonded to the transition metal, again confirming the distance-dependence of the enhancement effect. Comparison of the BHandH and SA-CASSCF NACs between the ground and second excited states shows that they are quite similar. This is not the case for the first and third excited state. In particular we notice that NACs between the ground and third excited state involving the N atoms are noticeably stronger with BHandH, which likely enables its contribution to the enhancement. We find that the non-adiabatic couplings computed with B3LYP closely match those obtained with BHandH, which accounts for the nearly identical enhanced VCD profiles predicted by both methods. NACs, difference plots and mean errors for other DFAs can be found in the SI.
(a) NACs between ground and first excited state
(b) NACs between ground and second excited state
(c) NACs between ground and third excited state